OpenStudy (anonymous):

A 50,000 kg locomotive is traveling at 10 m/s when its engine and brakes both fail. How far will the locomotive roll before it comes to a stop? Assume the track is level.

OpenStudy (doc.brown):

The train has no force accelerating it as the engine has failed. The only thing that will slow the train down are other forces, like air resistance, friction, and the x component of gravity (if the train starts up a hill). Without these forces, the train will roll forever. Do you have values for air resistance, or friction?

OpenStudy (anonymous):

no, but I would think that steel on steel coefficient of friction is available for everyone to look it up. http://www.tribology-abc.com/abc/cof.htm

OpenStudy (doc.brown):

\[F_{fr}=\mu_{k}F_{N}=\mu_{k}mg\]\[0.5(50000kg)(-9.8\frac{m}{s^{2}})\]\[=-245000N\] \[a=F/m\]\[-245000N/50000kg\]\[=-4.9\frac{m}{s^{2}}\]\[t = \frac{v-v_{0}}{a}\]\[\frac{0-10\frac{m}{s}}{-4.9\frac{m}{s^{2}}}\]\[=2s\]